OpenStudy (anonymous):

Solve A=πr+4t I dont know the answer :O

OpenStudy (anonymous):

me either

OpenStudy (saifoo.khan):

I go with e-shan

OpenStudy (anonymous):

until you tell me what I am supposed to do

OpenStudy (anonymous):

im sorry i forgot...

OpenStudy (anonymous):

Solve A=πr+4t for t.

OpenStudy (anonymous):

\[A - \pi*r = 4t\] \[\frac{A -\pi*r}{4}= t\]

OpenStudy (saifoo.khan):